# Adapt-Pruner: Adaptive Structural Pruning for Efficient Small Language Model Training

### v2

@@ 1 Introduction
Large language models (LLMs) (Kalyan, 2024; OpenAI, 2023) have demonstrated remarkable performance across a wide range of benchmarks. As their size increases, these models exhibit enhanced capabilities in understanding natural language and solving complex tasks through text generation (Zhao et al., 2023). However, achieving such performance requires models with billions of parameters, which presents significant challenges for practical deployment. The sheer scale of LLMs leads to high computational costs, making inference both resource-intensive and slow, and potentially introducing issues such as increased latency. Consequently, there is a growing demand for methods to compress LLMs (Zhu et al., 2024), aiming to reduce the number of parameters and improve inference speed, all while preserving the original model performance. Effective compression techniques hold the potential to create more efficient and deployable LLMs.
Several techniques have been proposed to compress LLMs, most of which fall into one of four categories: structured and unstructured pruning (Cheng et al., 2024), quantization (Gholami et al., 2022), low-rank factorization (Sainath et al., 2013), and knowledge distillation (Gou et al., 2021). In this paper, we primarily focus on structured pruning, which can be combined together with further training to obtain strong Small Language Models. Structured pruning removes entire filters, layers, or specific model components from neural networks, enabling both compression and realistic acceleration. On top of that, it does not require specialized hardware or library support to achieve these benefits (He and Xiao, 2023) in contrast to unstructured pruning.
@@
Furthermore, when computational resources allow, structural pruning methods can be combined together with additional post-training stages, recovering the performance hurt by pruning. Multiple approaches (Xia et al., 2024; Sreenivas et al., 2024) have investigated this fashion of training to obtain SLMs efficiently. Nonetheless, their methods fixate too much on the prune-the-train paradigm, which hinders them from further boosting the target SLM’s performance. In contrast, a different paradigm is explored in this paper, offering new opportunities for structural-pruning-based acceleration methods.
@@
A novel family of models called Adapt-LLMs, which is obtained through Adapt-Accel, achieves superior performance over strong open-sourced models. In particular, Adapt-Pruner recovers the performance of MobileLLM (Liu et al., 2024) in MMLU (Hendrycks et al., 2020) with less tokens via pruning from its larger counterparts. On top of that, a strong 1B model is discovered by pruning from DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI et al., 2025), outperforming Llama-3.2-1B (AI@Meta, 2024) in multiple benchmarks, including MMLU, TruthfulQA (Lin et al., 2021), and AGIEval (Zhong et al., 2023).
@@ 2 Related Work
Pruning removes weights and modifies the model’s architecture. Formally, given a neural network , pruning produces a new model , where is a binary mask that sets certain parameters to zero, and denotes element-wise multiplication. Pruning typically hurts the network’s performance, so post-training is often employed to recover the loss (Blalock et al., 2020). Overall, pruning methods can be categorized into two types, unstructured pruning and structured pruning. Unstructured pruning removes individual weights, resulting in sparsified weight matrices, but normally face difficulties in inference speedups when specialized hardware is unavailable (Dery et al., 2024). In contrast, structured pruning operates at a larger granularity, removing entire weight groups. This includes width pruning (Ma et al., 2023; Ashkboos et al., 2024), which removes groups of coupled weights, and depth pruning (Kim et al., 2024; Siddiqui et al., 2024), which eliminates entire layers. Our focus is on post-training structured pruning, balancing generality and hardware efficiency.
@@
Several works have explored adaptive compression. Dong et al. (2024) selects transformer feedforward experts and removes feedforward neurons during inference based on their high activation magnitudes from input prompts. While this method is effective, we seek an approach that can reduce model size without depending on specific input prompts. An et al. (2023) computes the sample variance of each input feature and weights it by the squared norm of the corresponding column in the weight matrix to determine a layer’s importance and assign sparsity accordingly. Wang and Kindratenko (2024) determines the optimal layer-wise sparsity distribution through reinforcement learning, where different sparsity patterns are sampled and updated based on their performance rewards. However, all aforementioned methods failed to take into account the loss on the functional aspect of models, where the overall mapping of the LLM is expected to be preserved during the pruning process to minimize the performance drop. Here and are input and output tensors of the model. Our method utilizes the mapping information to assign an adaptive sparsity across different decoder layers.
@@
Knowledge distillation is a technique used to transfer the advanced capabilities of high-performing LLMs to smaller models (Xu et al., 2024). Combining knowledge distillation with pruning can yield strong performance, where the original model acts as the teacher and the compressed model serves as the student (Sreenivas et al., 2024). Conventionally, knowledge distillation can be categorized into black-box (Ho et al., 2022; Hsieh et al., 2023) and white-box distillation (Gu et al., 2023; Latif et al., 2023; Agarwal et al., 2023; Zhou et al., 2023; Shum et al., 2024), depending on whether the model weights and prediction logits of the teacher model can be accessed. In that sense, structural-pruning-based acceleration methods can be roughly viewed as white-box knowledge distillation. However, in this paper, we focus on the adaptive pruning algorithm with supervised fine-tuning only. Though it is expected to achieve even better performance when integrated with state-of-the-art knowledge distillation techniques in post-training, it normally incurs more complexity and training cost, hence we leave that for future exploration.
@@
Training small language models from scratch demands substantial computational resources, making pruning of larger models with recovery post-training an attractive alternative. Recent works have explored various approaches to this challenge. Xia et al. (2024) proposes a systematic pruning framework that optimizes across four architectural dimensions (layer count, hidden dimension, attention heads, and MLP size) towards target architectures, followed by a continual post-training phase phase with dynamic batch loading. Sreenivas et al. (2024) leverages knowledge distillation, first fine-tuning the original model as a teacher and then transferring its knowledge to the pruned student model. However, these methods rely on single-phase post-training and do not exploit the potential benefits of incremental pruning with interleaved recovery phases, which is demonstrated to be effective in the experiments.
@@ 3.1.1 Assign Sparsity based on Importance
where denote the batch size, sequence length, and hidden size, respectively. Based on this, we use a function that measures the vector similarity or distance between and to assess the changes in the tensor caused by each decoder layer. The greater the similarity or the smaller the distance between and , the less important that decoder layer is. This intuition derives from Li et al. (2024), where it is observed that top self-attention layers have diminished gradient norms and those layers serve similar purposes as identity functions.
@@ 3.1.2 Pruning Weight Groups Inside Decoder Layer
To enable structure-aware pruning in Adapt-Pruner, methods from Ma et al. (2023); Fang et al. (2023) are employed to build dependency graphs for LLMs, which facilitates automatic identification and extraction of coupled structures in LLMs.
@@
With coupled structures and target sparsity defined for each group, the next step is selecting weight matrices to prune with minimized performance degradation. For any grouped weight structure , containing weight matrices, a calibration dataset is adopted to assess the relative importance of each matrix. Following (LeCun et al., 1989; Ma et al., 2023), the importance of the -th weight matrix in layer is defined as:
@@ 3.2 Adapt-Accel: Incremental Pruning with Interleaved Recovery Training
To achieve the target overall sparsity, our model undergoes multiple rounds of pruning, which inevitably leads to performance degradation. Inspired by neuroplasticity in biological neural networks (Walløe et al., 2014), periodic interleaved recovery phases are introduced through post-training after each pruning. The optimal pruning ratios for triggering these recovery phases were determined through ablation studies in Section 4.4.
4 Experiment
4.1 Adapt-Pruner as Effective LLM Pruners
@@
To demonstrate the effectiveness of Adapt-Pruner, different pruners are evaluated and compared on Llama-3.1-8B (AI@Meta, 2024). For task-agnostic performance evaluation of the pruned models, zero-shot classification is performaned on popular common-sense reasoning datasets: ARC-easy and ARC-challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019) and WinoGrande (Sakaguchi et al., 2021), where the averaged accuracy is reported. In particular, length-normalized accuracy is reported for any benchmarks that require length-dependent accuracy. Additionally, we supplement our evaluation with a generation task using WikiText2 (Merity et al., 2016). Following prior work (Ma et al., 2023; An et al., 2023; Ashkboos et al., 2024), we employ the LM Evaluation Harness (Gao et al., 2024) with default parameters, except that all models use the bfloat16 data type, and the batch size is set to ‘auto’ during evaluation.
@@ Results
As shown in Table 1, Adapt-Pruner outperforms all baselines by a non-trivial margin. In particular, Adapt-Pruner excels at preserving commonsense knowledge in LLMs during the pruning process. This is highly relevant to the skewed knowledge distribution across LLM layers, where the bottom layers are observed to be more important than the top layers (Pan et al., 2024; Li et al., 2024).
4.2 Adapt-Accel as Efficient LLM Trainers
Built on top of Adapt-Pruner, Adapt-Accel incorporates interleaved pruning and training to accelerate the optimization of SLMs, which is shown to be a better strategy compared to past methods (Xia et al., 2024; Sreenivas et al., 2024).
Setup
As the training of language models requires a non-trivial amount of computational resources, a smaller family of models are adopted for this section of experiments. MobileLLM (Liu et al., 2024) is a series of SLMs developed by Meta for on-device deployment purposes and stands for one of the strongest SLMs on the scale of 125M/350M.
To demonstrate the superiority of Adapt-Accel, three benchmarks are employed for evaluation, including BBH (Srivastava et al., 2022), TruthfulQA (Lin et al., 2021), AGIEval (Zhong et al., 2023), which assess all methods performance in different aspects beyond commonsense reasoning. A hybrid dataset with 3.87B tokens is utilized for training, where the data source is available in Table 2.
@@ 4.2 Adapt-Accel as Efficient LLM Trainers
| Model | Training Tokens | BBH (3-shot) | TruthfulQA (1-shot) | AGIEval (0-shot) | MMLU (5-shot) |
| MobileLLM-125M | 18.45 | 32.88 | 31.19 | 24.65 | |
| MobileLLM-350M 125M | 13.75 | 36.18 | 31.34 | 25.20 | |
| MobileLLM-350M | 19.98 | 29.86 | 30.85 | 26.11 | |
| MobileLLM-600M 350M | 23.61 | 34.22 | 31.93 | 32.50 | |
| MobileLLM-600M | 23.99 | 29.29 | 30.53 | 26.46 | |
| MobileLLM-1B 600M | 26.60 | 32.35 | 34.24 | 36.77 |
@@ Results
As shown in Table 4, Adapt-Accel outperforms both ShearedLLaMA (Xia et al., 2024) and NVIDIA’s Minitron Approach (Sreenivas et al., 2024) by a non-trivial margin. This demonstrates that the adaptively pruned models have preserved essential components in the original SLMs, and are still capable of learning new knowledge from the training set effectively.
| Model | Training Tokens | BBH (3-shot) | TruthfulQA (1-shot) | AGIEval (0-shot) | MMLU (5-shot) |
| MobileLLM-350M | 19.98 | 29.86 | 30.85 | 26.11 | |
| Qwen-2.5-0.5B 350M | 26.97 | 34.86 | 33.23 | 38.48 | |
| MobileLLM-1B | 25.88 | 30.26 | 31.34 | 24.98 | |
| LLaMA-3.2-1B | 31.01 | 30.41 | 31.68 | 31.27 | |
| Deepseek-R1-Distill-Qwen-1.5B 1B | 30.46 | 37.50 | 34.34 | 34.62 |
@@ 4.3 Adapt-LLMs as Strong LLMs
To provide evidence in support of the claimed strengths of Adapt-LLM, two types of experiments are conducted, individually demonstrating the acceleration and performance benefits of Adapt-LLM. The first branch of experiments focuses on the acceleration aspect of Adapt-Accel, where different sizes of MobileLLMs, ranging from 350M to 1B, are employed. The second branch of experiments emphasizes performance, where Adapt-Accel is applied to Qwen-2.5-0.5B (Team, 2024) and Deepseek-R1-Distill-Qwen-1.5B (DeepSeek-AI et al., 2025), proving that the pruned Adapt-LLMs can still match or even surpass popular strong open-source models without the heavy cost of pretraining. The same benchmarks and datasets of Section 4.2 are adopted here.
@@ Results
Table 5 further demonstrates the performance gain brought by Adapt-Accel via pruning from strong LLMs. It is worth noticing that Deepseek-R1-Distill-Qwen-1.5B (DeepSeek-AI et al., 2025) 1B leads to a 1B model even stronger than LLaMA-3.2-1B (AI@Meta, 2024) in MMLU, delivering three orders of magnitude cost reduction in terms of training tokens. This indicates that Adapt-Accel can serve as a favorable tool for inheriting performance from strong open-source LLMs, allowing researchers and engineers with limited computational resources to still keep up with the fast iteration speed of state-of-the-art LLMs.

### v3

@@ 1 Introduction
Large language models (LLMs) (KALYAN2024100048; openai2023gpt4) have demonstrated remarkable performance across a wide range of benchmarks. As their size increases, these models exhibit enhanced capabilities in understanding natural language and solving complex tasks through text generation (zhao2023survey). However, achieving such performance requires models with billions of parameters, which presents significant challenges for practical deployment. The sheer scale of LLMs leads to high computational costs, making inference both resource-intensive and slow, and potentially introducing issues such as increased latency. Consequently, there is a growing demand for methods to compress LLMs (zhu2024surveymodelcompressionlarge), aiming to reduce the number of parameters and improve inference speed, all while preserving the original model performance. Effective compression techniques hold the potential to create more efficient and deployable LLMs.
Several techniques have been proposed to compress LLMs, most of which fall into one of four categories: structured and unstructured pruning (cheng2024survey), quantization (gholami2022survey), low-rank factorization (sainath2013low), and knowledge distillation (gou2021knowledge). In this paper, we primarily focus on structured pruning, which can be combined together with further training to obtain strong Small Language Models. Structured pruning removes entire filters, layers, or specific model components from neural networks, enabling both compression and realistic acceleration. On top of that, it does not require specialized hardware or library support to achieve these benefits (he2023structured) in contrast to unstructured pruning.
@@
Furthermore, when computational resources allow, structural pruning methods can be combined together with additional post-training stages, recovering the performance hurt by pruning. Multiple approaches (xia2024shearedllamaacceleratinglanguage; sreenivas2024llmpruningdistillationpractice) have investigated this fashion of training to obtain SLMs efficiently. Nonetheless, their methods fixate too much on the prune-the-train paradigm, which hinders them from further boosting the target SLM’s performance. In contrast, a different paradigm is explored in this paper, offering new opportunities for structural-pruning-based acceleration methods.
@@
A novel acceleration paradigm called Adapt-Accel is presented, which is the first method that interleaves the pruning with training in a highly frequent manner and demonstrates non-trivial improvements compared to past methods (xia2024shearedllamaacceleratinglanguage; sreenivas2024llmpruningdistillationpractice).
@@
A novel family of models called Adapt-LLMs, which is obtained through Adapt-Accel, achieves superior performance over strong open-sourced models. In particular, Adapt-Pruner recovers the performance of MobileLLM (liu2024mobilellm) in MMLU (hendrycks2020mmlu) with less tokens via pruning from its larger counterparts. On top of that, a strong 1B model is discovered by pruning from DeepSeek-R1-Distill-Qwen-1.5B (deepseekr1), outperforming Llama-3.2-1B (llama3modelcard) in multiple benchmarks, including MMLU, TruthfulQA (lin2021truthfulqa), and AGIEval (zhong2023agieval).
@@ 2 Related Work
Pruning removes weights and modifies the model’s architecture. Formally, given a neural network , pruning produces a new model , where is a binary mask that sets certain parameters to zero, and denotes element-wise multiplication. Pruning typically hurts the network’s performance, so post-training is often employed to recover the loss (MLSYS2020_6c44dc73). Overall, pruning methods can be categorized into two types, unstructured pruning and structured pruning. Unstructured pruning removes individual weights, resulting in sparsified weight matrices, but normally face difficulties in inference speedups when specialized hardware is unavailable (dery2024everybodyprunenowstructured). In contrast, structured pruning operates at a larger granularity, removing entire weight groups. This includes width pruning (llm-pruner; SliceGPT), which removes groups of coupled weights, and depth pruning (kim2024shortenedllamadepthpruning; siddiqui2024deeperlookdepthpruning), which eliminates entire layers. Our focus is on post-training structured pruning, balancing generality and hardware efficiency.
@@
Several works have explored adaptive compression. dong2024prompt selects transformer feedforward experts and removes feedforward neurons during inference based on their high activation magnitudes from input prompts. While this method is effective, we seek an approach that can reduce model size without depending on specific input prompts. FLAP computes the sample variance of each input feature and weights it by the squared norm of the corresponding column in the weight matrix to determine a layer’s importance and assign sparsity accordingly. RLPruner determines the optimal layer-wise sparsity distribution through reinforcement learning, where different sparsity patterns are sampled and updated based on their performance rewards. However, all aforementioned methods failed to take into account the loss on the functional aspect of models, where the overall mapping of the LLM is expected to be preserved during the pruning process to minimize the performance drop. Here and are input and output tensors of the model. Our method utilizes the mapping information to assign an adaptive sparsity across different decoder layers.
@@
Knowledge distillation is a technique used to transfer the advanced capabilities of high-performing LLMs to smaller models (xu2024surveyknowledgedistillationlarge). Combining knowledge distillation with pruning can yield strong performance, where the original model acts as the teacher and the compressed model serves as the student (sreenivas2024llmpruningdistillationpractice). Conventionally, knowledge distillation can be categorized into black-box (ho2022large; hsieh2023distilling) and white-box distillation (gu2023knowledge; latif2023knowledge; agarwal2023gkd; zhou2023distillspec; shum2024first), depending on whether the model weights and prediction logits of the teacher model can be accessed. In that sense, structural-pruning-based acceleration methods can be roughly viewed as white-box knowledge distillation. However, in this paper, we focus on the adaptive pruning algorithm with supervised fine-tuning only. Though it is expected to achieve even better performance when integrated with state-of-the-art knowledge distillation techniques in post-training, it normally incurs more complexity and training cost, hence we leave that for future exploration.
@@
Training small language models from scratch demands substantial computational resources, making pruning of larger models with recovery post-training an attractive alternative. Recent works have explored various approaches to this challenge. xia2024shearedllamaacceleratinglanguage proposes a systematic pruning framework that optimizes across four architectural dimensions (layer count, hidden dimension, attention heads, and MLP size) towards target architectures, followed by a continual post-training phase phase with dynamic batch loading. sreenivas2024llmpruningdistillationpractice leverages knowledge distillation, first fine-tuning the original model as a teacher and then transferring its knowledge to the pruned student model. However, these methods rely on single-phase post-training and do not exploit the potential benefits of incremental pruning with interleaved recovery phases, which is demonstrated to be effective in the experiments.
@@ 3.1.1 Assign Sparsity based on Importance
where denote the batch size, sequence length, and hidden size, respectively. Based on this, we use a function that measures the vector similarity or distance between and to assess the changes in the tensor caused by each decoder layer. The greater the similarity or the smaller the distance between and , the less important that decoder layer is. This intuition derives from li2024mix, where it is observed that top self-attention layers have diminished gradient norms and those layers serve similar purposes as identity functions.
@@ 3.1.2 Pruning Weight Groups Inside Decoder Layer
To enable structure-aware pruning in Adapt-Pruner, methods from llm-pruner; depgraph are employed to build dependency graphs for LLMs, which facilitates automatic identification and extraction of coupled structures in LLMs.
@@
With coupled structures and target sparsity defined for each group, the next step is selecting weight matrices to prune with minimized performance degradation. For any grouped weight structure , containing weight matrices, a calibration dataset is adopted to assess the relative importance of each matrix. Following (lecun1989optimal; llm-pruner), the importance of the -th weight matrix in layer is defined as:
@@ 3.2 Adapt-Accel: Incremental Pruning with Interleaved Recovery Training
To achieve the target overall sparsity, our model undergoes multiple rounds of pruning, which inevitably leads to performance degradation. Inspired by neuroplasticity in biological neural networks (walloe2014stereological), periodic interleaved recovery phases are introduced through post-training after each pruning. The optimal pruning ratios for triggering these recovery phases were determined through ablation studies in Section 4.4.
4 Experiment
4.1 Adapt-Pruner as Effective LLM Pruners
@@
To demonstrate the effectiveness of Adapt-Pruner, different pruners are evaluated and compared on Llama-3.1-8B (llama3modelcard). For task-agnostic performance evaluation of the pruned models, zero-shot classification is performaned on popular common-sense reasoning datasets: ARC-easy and ARC-challenge (arc), HellaSwag (hellaswag), OpenBookQA (OpenBookQA), PIQA (piqa), SIQA (siqa) and WinoGrande (winogrande), where the averaged accuracy is reported. In particular, length-normalized accuracy is reported for any benchmarks that require length-dependent accuracy. Additionally, we supplement our evaluation with a generation task using WikiText2 (wikitext2). Following prior work (llm-pruner; FLAP; SliceGPT), we employ the LM Evaluation Harness (eval-harness) with default parameters, except that all models use the bfloat16 data type, and the batch size is set to ‘auto’ during evaluation.
@@ Results
As shown in Table 1, Adapt-Pruner outperforms all baselines by a non-trivial margin. In particular, Adapt-Pruner excels at preserving commonsense knowledge in LLMs during the pruning process. This is highly relevant to the skewed knowledge distribution across LLM layers, where the bottom layers are observed to be more important than the top layers (pan2024lisa; li2024mix).
@@
4.2 Adapt-Accel as Efficient LLM Trainers
Built on top of Adapt-Pruner, Adapt-Accel incorporates interleaved pruning and training to accelerate the optimization of SLMs, which is shown to be a better strategy compared to past methods (xia2024shearedllamaacceleratinglanguage; sreenivas2024llmpruningdistillationpractice).
Setup
As the training of language models requires a non-trivial amount of computational resources, a smaller family of models are adopted for this section of experiments. MobileLLM (liu2024mobilellm) is a series of SLMs developed by Meta for on-device deployment purposes and stands for one of the strongest SLMs on the scale of 125M/350M.
| Method |
12.29 | 34.88 | 30.53 | ||||||
12.56 | 34.13 | 30.36 | ||||||
| Adapt-Accel | 13.75 | 36.18 | 31.34 |
To demonstrate the superiority of Adapt-Accel, three benchmarks are employed for evaluation, including BBH (srivastava2022bbh), TruthfulQA (lin2021truthfulqa), AGIEval (zhong2023agieval), which assess all methods performance in different aspects beyond commonsense reasoning. A hybrid dataset with 3.87B tokens is utilized for training, where the data source is available in Table 2.
@@
As shown in Table 3, Adapt-Accel outperforms both ShearedLLaMA (xia2024shearedllamaacceleratinglanguage) and NVIDIA’s Minitron Approach (sreenivas2024llmpruningdistillationpractice) by a non-trivial margin. This demonstrates that the adaptively pruned models have preserved essential components in the original SLMs, and are still capable of learning new knowledge from the training set effectively.
@@ 4.3 Adapt-LLMs as Strong LLMs
To provide evidence in support of the claimed strengths of Adapt-LLM, two types of experiments are conducted, individually demonstrating the acceleration and performance benefits of Adapt-LLM. The first branch of experiments focuses on the acceleration aspect of Adapt-Accel, where different sizes of MobileLLMs, ranging from 350M to 1B, are employed. The second branch of experiments emphasizes performance, where Adapt-Accel is applied to Qwen-2.5-0.5B (qwen2.5) and Deepseek-R1-Distill-Qwen-1.5B (deepseekr1), proving that the pruned Adapt-LLMs can still match or even surpass popular strong open-source models without the heavy cost of pretraining. The same benchmarks and datasets of Section 4.2 are adopted here.
| Model | Training Tokens |
18.45 | 32.88 | 31.19 | 24.65 | |||||
13.75 | 36.18 | 31.34 | 25.20 | |||||
19.98 | 29.86 | 30.85 | 26.11 | |||||
23.61 | 34.22 | 31.93 | 32.50 | |||||
23.99 | 29.29 | 30.53 | 26.46 | |||||
26.60 | 32.35 | 34.24 | 36.77 |
@@ Results
Table 5 further demonstrates the performance gain brought by Adapt-Accel via pruning from strong LLMs. It is worth noticing that Deepseek-R1-Distill-Qwen-1.5B (deepseekr1) 1B leads to a 1B model even stronger than LLaMA-3.2-1B (llama3modelcard) in MMLU, delivering three orders of magnitude cost reduction in terms of training tokens. This indicates that Adapt-Accel can serve as a favorable tool for inheriting performance from strong open-source LLMs, allowing researchers and engineers with limited computational resources to still keep up with the fast iteration speed of state-of-the-art LLMs.
| Model | Training Tokens |
19.98 | 29.86 | 30.85 | 26.11 | |||||
26.97 | 34.86 | 33.23 | 38.48 | |||||
25.88 | 30.26 | 31.34 | 24.98 | |||||
31.01 | 30.41 | 31.68 | 31.27 | |||||
30.46 | 37.50 | 34.34 | 34.62 |